OpenStudy (anonymous):

In 1993, Mary was twice Ben's age. In 2000, she was 7 years older than Ben. How old is Ben in 2004?

OpenStudy (beginnersmind):

Use x to denote Ben's age in 1993. Can you express Mary's age in 1993 in terms of x? Mary's age in 2000? Ben's age in 2000?

OpenStudy (anonymous):

Mary in 1993 = 2x. 2000, 2x+7?

OpenStudy (beginnersmind):

Mary in 1993 = 2x. Correct 2000, 2x+7? Correct for Mary For Ben x+7 Now, can you express the sentence "In 2000, Mary was 7 years older than Ben" with an equation?

OpenStudy (beginnersmind):

If you're stuck, say so :)

OpenStudy (anonymous):

ahaha im stuck

OpenStudy (beginnersmind):

Ok, we've figured out that Ben's age in 1993: x Mary's age in 1993: 2*x (twice as much as Ben's) Ben's age in 2000: x+7 Mary's age in 2000: 2x+7 We know that in 2000 Mary was 7 years older than Ben. So if we add 7 to Ben's age in 2000, we get Mary's age in 2000. We can write it as (x+7)+7=2*x+7 Do you follow so far? If not, which step is unclear?

OpenStudy (anonymous):

Oh yep

OpenStudy (anonymous):

@beginnersmind

OpenStudy (beginnersmind):

Can you solve the last equation: (x+7)+7=2x+7 ?

OpenStudy (anonymous):

aha oops x=7

OpenStudy (anonymous):

Thanks :D

OpenStudy (beginnersmind):

Ok, remember that x was Ben's age in 1993. They asked for Ben's age in 2004

OpenStudy (anonymous):

In 2004 he will be 18 (:

OpenStudy (beginnersmind):

Right :) How do you do that inverted smiley-face?

OpenStudy (anonymous):

LOL all you do is press the brackets first (:

OpenStudy (beginnersmind):

LOL, I feel silly now

OpenStudy (anonymous):

AHA it's alright